OpenStudy (anonymous):

how do i prove this: sin2(x) (cot(x) + tan (x)) =2

OpenStudy (anonymous):

rewrite everything with sine and cosine then multiply

OpenStudy (anonymous):

you can put \(a=\cos(x), b=\sin(x)\) and start with \[b^2(\frac{a}{b}+\frac{b}{a})\]

OpenStudy (anonymous):

\[2\sin \theta \cos \theta (\frac{ \cos }{ \sin }\left(\begin{matrix}\sin \\ \cos\end{matrix}\right) )\] have this so far

OpenStudy (anonymous):

would that make it \[\cos^2 x \sin^2 \over (\sin)(\cos)\]

OpenStudy (anonymous):

ok now i am lost are you trying to solve that for x? or are you trying to prove it is always equal to two?

OpenStudy (anonymous):

trying to prove it's equal to 2

OpenStudy (anonymous):

it isn't so don't try